# LOTFormer: Doubly-Stochastic Linear Attention via Low-Rank Optimal Transport

### v1

@@ 4.2 Long Range Arena
We evaluate our approach on the Long Range Arena (LRA) benchmark, a widely used suite for testing the ability of sequence models to capture long-range dependencies across diverse modalities, including structured parsing (ListOps), natural language understanding (Text), image classification (Image), document retrieval (Retrieval), and synthetic reasoning (Pathfinder). Following PolaFormer (Meng et al., 2025) and FlattenAttention (Han et al., 2023), we optionally apply a channel-wise 1D depthwise convolution over the token dimension to the value stream prior to mixing:
$\tilde{V}=\mathrm{DWConv}_{k}(V),\qquad O=A\,\tilde{V}$,
where $A$ is the attention map (DS for non-[CLS] rows; [CLS] via its head). DWC injects a local inductive bias and improves token mixing at negligible parameter/FLOP cost.
Results. We evaluate LOTFormer on LRA with $r=32$ both with and without DWC and present our results in Table 1. It can be seen that, without using a DWC, our method achieves the second highest average performance, beaten only by Polaformer Meng et al. (2025) which utilizes a DWC. By adding the DWC, we significantly boost our performance on both the text and image tasks, achieving the best average performance among all methods while also achieving the best performance on the image task and second highest performance on the text task.
Model Text ListOps Retrieval Pathfinder Image Avg.
@@
PolaFormerα=3 73.06 37.35 80.50 70.53 42.15 60.72
LOTFormer 65.2 $\pm$ 0.2 38.5 $\pm$ 1.2 80.4 $\pm$ 0.5 73.2 $\pm$ 0.9 45.7 $\pm$ 1.2 60.6
+ DWC 71.1 $\pm$ 1.0 38.5 $\pm$ 1.4 80.9 $\pm$ 0.2 69.9 $\pm$ 0.5 54.1 $\mathbf{\pm}$ 1.5 62.9
Table 1: Results on the LRA benchmark. The LOTFormer results are averaged over three runs. Best numbers per column are in bold. Our method achieves the highest overall average accuracy, with particularly strong gains on the Image task.
4.3 ImageNet-1K
Next, we evaluate LOTFormer on ImageNet-1K with a DeiT-Tiny (Touvron et al., 2021) backbone and compare against recent linear-attention variants (Table 2). When applying doubly stochastic attention in ViTs, special care is needed for the [CLS] token, since it must act as a hub that aggregates information from all other tokens; a naive application may restrict its ability to gather global context. In addition, many recent linear-attention variants achieve improvements through orthogonal techniques that are compatible with LOTFormer. For example, PolaFormer and FlattenAttention employ lightweight depthwise convolutions (DWC) on values, while PolaFormer also introduces polarization on the logits— both of which could be seamlessly incorporated into our framework. In this section, we first discuss our special treatment of the [CLS] token, and then provide a thorough study on ImageNet-1K to disentangle the effects of these complementary techniques and our proposed LOTFormer. In short, we demonstrate that LOTFormer with a [CLS]-specific polarization head (Pol-[CLS]) and a lightweight DWC on values reaches 74.8% at negligible overhead (see ablations in Table 3), surpassing strong baselines such as PolaFormer (74.6%) at the same resolution.
[CLS] under DS attention.
In the extreme case (no entropy regularization and $r=n$), a doubly stochastic attention matrix can degenerate into a permutation, causing each query to attend to only one key—so the [CLS] token would aggregate values from just a single token. Although entropy regularization and low-rank constraints mitigate this, the issue persists in principle and conflicts with the [CLS] token’s role as a global aggregator. To address this, we retain standard softmax aggregation for the [CLS] token, while enforcing doubly stochasticity only among image tokens. This preserves the classic softmax attention for [CLS]:
$\alpha_{\mathrm{cls},j}=\frac{\exp\!\big(\beta\,\langle q_{\mathrm{cls}},k_{j}\rangle\big)}{\sum_{l}\exp\!\big(\beta\,\langle q_{\mathrm{cls}},k_{l}\rangle\big)},\qquad\beta>0,$
which decouples [CLS] from the DS row–column coupling and restores true global aggregation, while all non-[CLS] rows remain DS.
Polarization variants: Pol-[CLS] vs. Pol-All.
Following Meng et al. (2025), let $x^{+}=\mathrm{ReLU}(x)$ and $x^{-}=\mathrm{ReLU}(-x)$. Pol-[CLS] (ours for LOTFormer) sharpens only the [CLS] interactions by forming polarized logits
$\tilde{s}_{\mathrm{cls},j}=\big((q_{\mathrm{cls}}^{+})^{\top}k_{j}^{+}+(q_{\mathrm{cls}}^{-})^{\top}k_{j}^{-}\big)^{p_{s}}+\big((q_{\mathrm{cls}}^{+})^{\top}k_{j}^{-}+(q_{\mathrm{cls}}^{-})^{\top}k_{j}^{+}\big)^{p_{o}},$
followed by softmax over $j$: $\alpha_{\mathrm{cls},j}=\mathrm{softmax}_{j}(\tilde{s}_{\mathrm{cls},j})$. By contrast, Pol-All (as in PolaFormer) applies the same polarized-logit construction to every row $i$, then normalizes each row with softmax: $\alpha_{i,j}=\mathrm{softmax}_{j}(\tilde{s}_{i,j})$. Non-[CLS] rows in LOTFormer remain DS; only the [CLS] row uses softmax (or Pol-[CLS]+softmax). We reflect this choice in tables as Head/Pol $\in\{\text{Softmax},\ \text{Pol-{[CLS]} (softmax)},\ \text{Pol-All (softmax)}\}$.
Depthwise convolution on values (DWC).
We study DWC in both PolaFormer (Pol-All) and LOTFormer (Pol-[CLS]) and observe consistent gains; our best LOTFormer uses Pol-[CLS] (softmax) + DWC.
Reporting protocol.
Table 2 reports the best configuration per method at the same input resolution and discloses two implementation choices: the [CLS] Head/Pol (Softmax / Pol-[CLS] (softmax) / Pol-All (softmax)) and whether DWC is used. DS methods (Sinkformer, ESPFormer, LOTFormer) adopt [CLS]–softmax to enable global aggregation; only LOTFormer evaluates Pol-[CLS]. PolaFormer uses Pol-All and DWC. Detailed ablations isolate the effects of (i) the [CLS] head and (ii) DWC.
Table 2: DeiT-Tiny on ImageNet-1K with linear-attention variants. Best in bold; other top-3 are underlined. DS methods use [CLS]–softmax for the [CLS] row; Pol-* entries use softmax kernels (polarized logits + softmax).
Method Reso Params FLOPs [CLS] Head/Pol DWC Top-1 (%)
DeiT (Touvron et al., 2021) $224^{2}$ 5.7M 1.1G Softmax 72.2
EfficientAttn (Shen et al., 2021) $224^{2}$ 5.7M 1.1G Softmax 70.2
HydraAttn (Bolya et al., 2022) $224^{2}$ 5.7M 1.1G Softmax 68.3
EnhancedAttn (Cai et al., 2022) $224^{2}$ 5.8M 1.1G Softmax 72.9
FLattenAttn (Han et al., 2023a) $224^{2}$ 6.1M 1.1G Softmax 74.1
AngularAttn (You et al., 2023) $224^{2}$ 5.7M 1.1G Softmax 70.8
MobiAttn (Yao et al., 2024) $224^{2}$ 5.7M 1.2G Softmax 73.3
PolaFormer $224^{2}$ 6.1M 1.2G Pol-All (softmax) 74.6
Sinkformer $224^{2}$ 5.7M 1.2G Softmax 70.2
ESPFormer $(L{=}8,\tau{=}0.1)$ $224^{2}$ 5.7M 1.2G Softmax 70.1
@@ 4.3 ImageNet-1K
Effect of $r$.
Accuracy improves with larger pivot size $r$ and saturates around r=32r{=}326464. The Pol-[CLS] + DWC variant consistently outperforms the base DS configuration at each $r$, indicating that a specialized [CLS] aggregator and lightweight local mixing are complementary to the low-rank transport factorization. We use r=32r{=}32 by default for DeiT-Tiny as it offers a strong accuracy–efficiency trade-off.
@@
4.4 Neural Machine Translation
We evaluate LOTFormer against two recent attention mechanisms that also rely on doubly stochastic matrices, namely Sinkformer and ESPFormer. For fairness, we embed all three variants into two standard reference backbones: the Transformer and its DiffTransformer counterpart Ye et al. (2025). Both backbones are implemented using the fairseq sequence modeling toolkit Ott et al. (2019) and trained on the IWSLT’14 German-to-English dataset Cettolo et al. (2014). Each backbone consists of a 6-layer encoder and a 6-layer decoder.
We adopt a two-stage evaluation protocol. In the first stage, we pre-train the baseline Transformer and DiffTransformer models for 25 epochs using the standard training procedure. We then conduct a Plug-and-Play evaluation, where the attention heads of LOTFormer, ESPFormer, and Sinkformer are directly inserted into the pre-trained models and evaluated without further training. The results of this evaluation are summarized in Table 5.
In the second stage, we perform a Fine-Tune Boost phase, where the Plug-and-Play models are further fine-tuned for an additional 10 epochs. Fine-tuning consistently improves performance across all three doubly stochastic variants. Notably, LOTFormer achieves the largest gains, surpassing both ESPFormer and Sinkformer on the Transformer and DiffTransformer backbones, and attaining the best BLEU scores of 34.64 and 34.83, respectively. These results demonstrate the effectiveness of LOTFormer as a transport-based doubly stochastic attention mechanism.
Table 5: Plug-and-Play and Fine-Tune performance on IWSLT’14 De→En (median over 4 runs). Note that indicates plug-and-play performance when swapping in a different attention than the base model’s own. Moreover, $\Delta$ shows the change vs. the base model in each block.
Base: Transformer Base: DiffTransformer
Model Plug-and-Play Fine-Tune $\Delta$ Plug-and-Play Fine-Tune $\Delta$
Transformer 33.40 34.61 33.85 34.78
Sinkformer 33.36 34.61 +0.00 33.67 34.81 +0.03
ESPFormer 33.38 34.64 +0.03 33.72 34.83 +0.05
LOTFormer (ours) 33.29 34.72 +0.11 33.42 34.91 +0.13
4.5 Ablation Study
Table 6: Ablation results for LOTFormer. Top-1 Accuracy (%) on ImageNet-100.
(i) Vary $T$ @ $\varepsilon\in\{0.1,1\}$
(ii) Vary $\varepsilon$ @ T=10T{=}10
$0.01$ $0.1$ 11 $10$ $100$ fixed $\mathsf{ref}$ $79.1$ $80.8$ $81.8$ $81.2$ $79.5$ learnable $\mathsf{ref}$ $79.2$ $81.1$ 82.1 $81.3$ $79.6$
Here we ablate different design choices in our experiments and present them in Table 6. All ablation studies were carried out on ImageNet-100, and with a down-sized DeiT Tiny with halved hidden dimensions. We first compare a fixed vs. learnable $\mathsf{ref}$ with uniform mass. At tighter entropies ($\varepsilon{=}0.1$), performance improves steadily as $T$ increases, e.g., fixed: $79.6\!\to\!80.9$ (from T=1T{=}1 to $20$), learnable: $79.8\!\to\!81.1$, reflecting that tighter $\varepsilon$ requires more Sinkhorn iterations to converge. At a looser setting ($\varepsilon{=}1$), both variants climb more quickly and saturate earlier (fixed: $80.2\!\to\!81.8$ at T=10T{=}10, learnable: $80.5\!\to\!\mathbf{82.1}$ at T=10T{=}10), with the learnable $\mathsf{ref}$ consistently ahead by $\approx\!0.3$$0.4$ points. The bottom panel (varying $\varepsilon$ at T=10T{=}10) shows the same pattern from the complementary view: accuracy peaks near $\varepsilon{=}1$ (fixed $81.8$, learnable $\mathbf{82.1}$) and drops at extremes (e.g., $\varepsilon{=}0.01$ or $100$), indicating under-converged or over-smoothed transport when $T$ is held fixed. Overall, a learnable $\mathsf{ref}$ matches or exceeds a fixed one across settings and reaches the best value of $\mathbf{82.1}\%$ at $(\varepsilon{=}1,\,T{=}10)$.
@@ 5 Conclusion
[…]
Acknowledgments
SK acknowledges support from the NSF CAREER Award No. 2339898 and the Wellcome Leap “Surgery: Assess/Validate/Expand (SAVE)” program. KH was partially supported by a Research Enhancement Program grant from the College of Science at the University of Texas at Arlington and by the Army Research Office under Grant W911NF-23-1-0213. The views and conclusions contained herein are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Notations and conventions
In the discrete measure setting, i.e., $\mu=\sum_{i=1}^{n}p^{1}_{i}\delta_{x_{i}},\nu=\sum_{j=1}^{m}p^{2}_{j}\delta_{y_{j}}$, given $r\leq n,m$, the (entropic) low rank optimal transport problem Scetbon et al. (2021); Scetbon & Cuturi (2022) is defined as
$LrOT_{r}(\mu,\nu):=\min_{\gamma\in U(\mathrm{p}^{1},\mathrm{p}^{2};r)}\sum_{i,j}c(x_{i},y_{j})\Gamma_{i,j}-\epsilon H(\Gamma).$ (B.1)
@@
From Theorem 3.2 in Cohen & Rothblum (1993) (Also see (5) in Scetbon & Cuturi (2022)), we have
@@
Thus, the low-rank optimal transport (B.1) becomes
$LrOT_{r}(\mu,\nu)$ $=\min_{\mathrm{p}^{0}\in\Delta_{r}^{+}}\min_{\begin{subarray}{c}\Gamma=(\Gamma^{1})^{\top}\text{diag}(1/\sigma)\Gamma^{2}:\\ \Gamma^{1}\in U(\mathrm{p}^{0},\mathrm{p}^{1}),\Gamma^{2}\in U(\mathrm{p}^{0},\mathrm{p}^{2})\end{subarray}}\sum_{i,j}c(x_{i},y_{j})\Gamma_{i,j}-\epsilon H(\Gamma)$ (B.2)
@@
$LOT(\mu,\nu;\sigma)=\sum_{i,j}c(x_{i},y_{j})\Gamma_{i,j},\Gamma=(\Gamma^{1})^{\top}\text{diag}(1/\sigma)\Gamma^{2}+\epsilon(\Gamma)$ (B.3)
@@
$LOT_{r}(\mu,\nu):=\inf_{\sigma\in\mathcal{P}_{r}(\mathbb{R}^{D})}LOT(\mu,\nu;\sigma)$ (B.4)
@@
For each $\sigma$, from the definition of $LOT(\mu,\nu;\sigma)$, we the $\gamma^{1},\gamma^{2}$ in the definition of LOT (B.3) satisfy $\gamma^{1}\in\Gamma(\mathrm{p}^{0},\mathrm{p}^{1}),\gamma^{1}\in\Gamma(\mathrm{p}^{1},\mathrm{p}^{2})$ and $\mathrm{p}^{0}\in\Delta_{r}^{+}$ by the definition of $\sigma$. Thus,

### v2

@@ 4.3 Long Range Arena
We evaluate LOTFormer on the Long Range Arena (LRA) benchmark, which stresses long context modeling across five modalities: structured parsing (ListOps), text classification (Text), image classification (Image), document retrieval (Retrieval), and synthetic reasoning (Pathfinder). To make the comparison as controlled as possible, we follow the same LRA training and preprocessing recipe used by Skyformer, including the two layer Transformer architecture and matching optimization settings.
Results and comparison to linear attention and state space models. Table 2 summarizes LRA performance for LOTFormer and a broad set of efficient baselines. We focus on recent linear time attention methods and modern state space models, and include strong SSM references such as DeltaNet (Yang et al., 2025) and Mamba2 (Dao and Gu, 2024), which are competitive alternatives on long sequences. Following common LRA recipes, we report LOTFormer both in the plain Transformer form and with an DWC token mixing layer, since several established baselines rely on this component as part of their model design. For a direct and controlled comparison, we adopt the same Transformer configuration and optimization choices used in Skyformer(Chen et al., 2021). Across tasks, LOTFormer matches or improves upon prior baselines under the same protocol and remains competitive with strong SSM alternatives. When DWC is enabled, LOTFormer attains its best average performance, indicating that low rank doubly stochastic normalization provides a practical linear time route to stable long context modeling on LRA.
Table 2: LRA results. LOTFormer numbers are means over three runs. Best per column in bold. $\dagger$ indicates +DWC.
Model Text List Ret Path Img Avg
@@
PolaFormerα=3 59.50 38.80 79.50 67.50 42.00 57.46
@@
Table 4: Plug-and-Play conversion on fine-tuned GLUE using BERT-base. We replace encoder bidirectional self-attention blocks only while keeping other parameters unchanged. Percent recovery is computed relative to the fine-tuned baseline.
@@
CoLA 58.8 43.6 56.9 59.2 59.1
SST2 93.2 87.7 90.9 92.6 92.5
MRPC 90.2 83.0 89.1 90.1 90.3
STS–B 88.8 78.6 77.7 87.4 89.7
QQP 91.0 86.7 90.0 91.0 90.5
MNLI 84.7 78.9 77.4 82.6 83.0
QNLI 91.3 84.6 84.5 89.6 90.2
RTE 68.2 54.1 56.3 69.3 68.5
Recover (%) 100.0 88.9 93.5 99.3 99.7
4.4 Plug-and-Play on Text benchmarks
A central goal of this work is to make doubly stochastic attention usable as a drop-in replacement in existing Transformer checkpoints. To this end, our Plug-and-Play evaluation starts from a trained model and performs a surgical conversion: we replace only the encoder bidirectional self-attention operator with a candidate attention mechanism, while keeping the rest of the network bitwise unchanged (token embeddings, FFNs, layer norms, residual connections, positional encodings, and task-specific heads). This protocol treats attention as a modular component and isolates its contribution without confounding changes from retraining. For encoder–decoder models, we leave the decoder causal self-attention and encoder–decoder cross-attention untouched, and convert only the encoder self-attention blocks.
[…]
GLUE conversion on fine-tuned BERT. For GLUE (Wang et al., 2019), we follow the same fine-tuned conversion protocol used in Hedgehog (Zhang et al., 2024a) and Transformer-to-RNN (T2R) (Kasai et al., 2021): starting from a task fine-tuned BERT-base (Devlin et al., 2019) checkpoint, we swap all encoder bidirectional self-attention blocks with LOTFormer, run a short conversion stage using an distillation objective, and then fine-tune on the original task to measure recovery relative to softmax.
[…]
Model Plug and Play Fine Tune $\Delta$
Base: Transformer
[…]
Sinkformer 33.36 34.61 +0.00
ESPFormer 33.38 34.64 +0.03
[…]
Base: DiffTransformer
[…]
Sinkformer 33.67 34.81 +0.03
ESPFormer 33.72 34.83 +0.05
[…]
4.5 Ablation Studies
Tables 6 and 7 study key design choices for LOTFormer under two representative settings.
Sinkhorn settings and the reference measure on ViT. Table 6 varies the number of Sinkhorn iterations $T$, the entropic strength $\varepsilon$, and the choice of reference measure $\mathsf{ref}$, either fixed or learnable and initialized to uniform. Experiments use ImageNet 100 with a reduced ViT backbone that has half the hidden dimension. For tighter regularization with $\varepsilon{=}0.1$, accuracy improves as $T$ increases, which is consistent with needing more iterations to approach a well balanced coupling. For $\varepsilon{=}1$, performance saturates earlier and the learnable $\mathsf{ref}$ is consistently stronger, reaching its best value at T=10T{=}10. Sweeping $\varepsilon$ at fixed T=10T{=}10 shows a clear optimum near $\varepsilon{=}1$, with degraded accuracy at both extremes. Very small $\varepsilon$ is harder to optimize under a fixed iteration budget, while very large $\varepsilon$ over smooths the transport plan. Overall, making $\mathsf{ref}$ learnable matches or improves on a fixed reference across all settings, and we use it in our strongest configurations.
Pivot size $r$ on DeiT Tiny. Table 7 quantifies the accuracy and cost trade off as we vary the pivot size $r$ for a strong DeiT Tiny configuration on ImageNet 1K. Increasing $r$ improves Top 1 up to a point, peaking at r=32r{=}32 and then slightly decreasing at r=64r{=}64. As expected from the $O(nr)$ structure, both forward and backward time increase monotonically with $r$. This underscores that $r$ is a single knob controlling the accuracy–efficiency tradeoff, motivating per backbone tuning of $r$ under a fixed compute budget.
Table 6: Ablations for LOTFormer on a ViT backbone. Top 1 accuracy (%) on ImageNet 100.
(i) Vary $T$ at $\varepsilon\in\{0.1,1\}$
(ii) Vary $\varepsilon$ at T=10T{=}10
$\varepsilon$=$0.01$ $\varepsilon$=$0.1$ $\varepsilon$=11 $\varepsilon$=$10$ $\varepsilon$=$100$ Fixed $\mathsf{ref}$ $79.1$ $80.8$ $81.4$ $81.2$ $79.5$ Learnable $\mathsf{ref}$ $79.2$ $81.1$ 82.1 $81.3$ $79.6$
@@ 5 Conclusion
We introduced LOTFormer, a transport based attention mechanism that is both linear time and doubly stochastic. By conditioning on a learnable low rank pivot measure, LOTFormer composes two entropic OT couplings, to induce a rank $r$ attention operator that applies to values in $O(nr)$ time without ever forming an $n\times n$ attention matrix. Across settings, LOTFormer delivers strong empirical results: it is competitive on ImageNet 1K under standard training recipes, it enables direct plug and play replacement of softmax attention in pretrained checkpoints by swapping only the attention blocks, and it achieves superior performance on Long Range Arena. Overall, these results show that transport structured normalization can be practical at scale, providing an efficient path to doubly stochastic attention without sacrificing accuracy.
Impact Statement
This work proposes LOTFormer, a linear time doubly stochastic attention operator derived from an optimal transport factorization that avoids forming an $n\times n$ attention map and is designed to be used as a drop in module. Its main intended impact is practical: under a Plug and Play protocol we can convert trained checkpoints by swapping only encoder bidirectional self attention while keeping the rest of the network fixed, enabling cheaper and more controlled comparisons of attention mechanisms.The approach targets bidirectional encoder attention, since standard causal masking makes doubly stochastic constraints degenerate. As with most efficiency improvements, lower cost can reduce per run compute but can also increase overall usage, and it may indirectly enable misuse, so deployment should follow standard dataset and application safety practices.
@@ Appendix A Notations and conventions
In the discrete measure setting, i.e., $\mu=\sum_{i=1}^{n}p^{1}_{i}\delta_{x_{i}},\nu=\sum_{j=1}^{m}p^{2}_{j}\delta_{y_{j}}$, given $r\leq n,m$, the (entropic) low rank optimal transport problem (Scetbon et al., 2021; Scetbon and Cuturi, 2022) is defined as
$LrOT_{r}(\mu,\nu):=\min_{\gamma\in U(\mathrm{p}^{1},\mathrm{p}^{2};r)}\sum_{i,j}c(x_{i},y_{j})\Gamma_{i,j}-\epsilon H(\Gamma).$ (7)
@@
From Theorem 3.2 in (Cohen and Rothblum, 1993) (Also see (5) in (Scetbon and Cuturi, 2022)), we have
@@
Thus, the low-rank optimal transport (7) becomes
$LrOT_{r}(\mu,\nu)$ $=\min_{\mathrm{p}^{0}\in\Delta_{r}^{+}}\min_{\begin{subarray}{c}\Gamma=(\Gamma^{1})^{\top}\text{diag}(1/\sigma)\Gamma^{2}:\\ \Gamma^{1}\in U(\mathrm{p}^{0},\mathrm{p}^{1}),\Gamma^{2}\in U(\mathrm{p}^{0},\mathrm{p}^{2})\end{subarray}}\sum_{i,j}c(x_{i},y_{j})\Gamma_{i,j}-\epsilon H(\Gamma)$ (8)
@@
$LOT(\mu,\nu;\sigma)=\sum_{i,j}c(x_{i},y_{j})\Gamma_{i,j},\Gamma=(\Gamma^{1})^{\top}\text{diag}(1/\sigma)\Gamma^{2}+\epsilon(\Gamma)$ (9)
@@
$LOT_{r}(\mu,\nu):=\inf_{\sigma\in\mathcal{P}_{r}(\mathbb{R}^{D})}LOT(\mu,\nu;\sigma)$ (10)
@@
For each $\sigma$, from the definition of $LOT(\mu,\nu;\sigma)$, we the $\gamma^{1},\gamma^{2}$ in the definition of LOT (9) satisfy $\gamma^{1}\in\Gamma(\mathrm{p}^{0},\mathrm{p}^{1}),\gamma^{1}\in\Gamma(\mathrm{p}^{1},\mathrm{p}^{2})$ and $\mathrm{p}^{0}\in\Delta_{r}^{+}$ by the definition of $\sigma$. Thus,